OpenStudy (praetorian.10):

2 questions for 2 manly cuddles xD

OpenStudy (praetorian.10):

Michigan's Isle Royale is located in Lake Superior and is 45 miles long and 9 miles wide. Ecologists estimate that the moose population has been on the island since around 1900. The wolf population, on the other hand, has only been there since the winter of 1949. The conditions had to be just right for the lake to freeze and the weather to be calm enough for the first wolves to make the passage over to the island. When the wolves arrived, they were greeted with a large moose population that had no predators. The moose population was at a dangerously high level, but was kept in check by the arrival of the wolves. This "harmonious" pattern continued for at least 20 years.

OpenStudy (praetorian.10):

In 1982, the wolf population suddenly crashed. The problem was a new canine virus. This virus surfaced around the world in 1978 and spread rapidly in about six months. Infected animals can survive, but they have to be treated immediately. Ecologists hypothesized that the virus arrived on the island carried on the shoe of a human traveller. The wolf population decreased from 50 to 14 individuals in a year. While the wolf population dropped, the moose population increased. However, the winter of 1996 was extremely harsh and the moose population experienced a major crash. More than 500 animals died during the winter months, reducing the population to less than 800 animals.

OpenStudy (praetorian.10):

What most likely happened when the wolves first arrived on the island? (A) The wolf population was at its carrying capacity. (B) The wolf population increased exponentially. (C) The moose population crashed because of the introduction of a predator. (D) The wolf population decreased because of the introduction of a virus. (E) The moose population increased exponentially.

OpenStudy (praetorian.10):

During a one-year study, a graduate student attempted to count the wolves and moose on the island. He estimated that there were 950 moose and 120 wolves. How many organisms, populations, and species were included in this study? (A) Organisms = 1,070; populations = 1; species = 2. (B) Organisms = 2; populations = 1; species = 2. (C) Organisms = 1,070; populations = 2; species = 2. (D) Organisms = 2; populations = 2; species = 1,070. (E) None of the above.

OpenStudy (praetorian.10):



OpenStudy (anonymous):

(D) The wolf population decreased because of the introduction of a virus. (D) Organisms = 2; populations = 2; species = 1,070.

OpenStudy (leonardo0430):

yes @endie4619 both answers are D. :)..................................i demand for my cuddle and cookies naow! ^_^

OpenStudy (anonymous):

i knew that

OpenStudy (praetorian.10):